Consider 2 rectangles: 3 by 8 and 4 by 8. Find the Area and Perimeter for each rectangle. Use your work to deduce that it is possible for a (non-square) rectangle's Area and Perimeter to be (numerically) equal. Can you find such a rectangle where one side is length 8?
Found 2 solutions by ikleyn, ewatrrr:
Answer by ikleyn(52817)

.

Good question for an advanced alumni of a kindergarten.


Do not understand, what is the need to post it to the forum.


Can you explain it, PLEASE ?



Answer by ewatrrr(24785)


Hi
A - wh  and P = 2w + 2L
3x8  A = 24  P = 22  numerically
4x8  A = 32  P = 24  numerically
(non-square) rectangle's Area and Perimeter to be (numerically) equal. 
Can you find such a rectangle where one side is length 8?
 8w = 16 + 2w  numerically
  6w = 16
   w = 8/3 
(8/3)x 8

64/3 = 48/3 + 16/3  checks.
Wish You the Best in your Studies.